# Direct Preference Optimization of Video Large Multimodal Models from Language Model Reward

### v1

@@ 4 Assessment of Evaluator with GPT-4V Caption as Evidence
FrozenBiLM (Yang et al., 2022)*** 1B 32.2 - 16.8 - 41.0 -
VideoLLaMA (Zhang et al., 2023a)*** 7B 51.6 2.5 29.6 1.8 - -
LLaMA-Adapter (Zhang et al., 2023b)*** 7B 54.9 3.1 43.8 2.7 - -
VideoChat (Li et al., 2023b)*** 7B 56.3 2.8 45.0 2.5 34.4 2.3
BT-Adapter Liu et al. (2023c)*** 7B 67.5 3.7 57.0 3.2 - -
@@
Table 1: Evaluation of Model Performance on Zero-Shot Video Question Answering Benchmarks Using gpt-3.5-turbo-0613. Models denoted with *** have their results directly sourced from their original publications. Caution is advised when interpreting these results; see Appendix A for an in-depth analysis of evaluation challenges. All other baseline models were reproduced by our team.
@@ Appendix F Prompts for GPT-4V and ChatGPT Queries
Task Instructions:
[…]
%**** 90_appendix.tex Line 50 ****{caption}
[…]
Q1: <question1>
A1: <answer1>
Q2: <question2>
A2: <answer2>
Q3: <question3>
A3: <answer3>
@@
Your role is to serve as an impartial and objective evaluator of a video caption provided by a Large Multimodal Model (LMM). Based on the input frames of a video, assess primarily on two criteria: the coverage of video elements in the caption and the absence of hallucinations in the response. In this context, hallucination refers to the model generating content not present or implied in the video, such as incorrect details about objects, actions, counts, or other aspects not evidenced in the video frames.
[…]
Then, assign a rating from the following scale:
[…]
Rating 5: Very informative, no hallucination
Rating 4: Somewhat informative with some missing details, no hallucination
Rating 3: Not informative, no hallucination
Rating 2: Very informative, with hallucination
Rating 1: Somewhat informative, with hallucination
Rating 0: Not informative, with hallucination
[…]
{LLM_response}
[…]
Judgment: <your judgment>
Score: <integer value rating>
@@
Given the following inputs:
[…]
2. **Question Related to the Caption**: {question}
3. **Ground Truth Answer**: {answer}
4. **Model Predicted Answer**: {prediction}
[…]
- **Accuracy**: Compare the predicted answer to the ground truth answer. Does the prediction accurately reflect the information given in the ground truth answer without introducing factual inaccuracies?
- **Clarity**: Assess the clarity of the predicted answer. Look for issues such as repetition, unclear descriptions, or any grammatical errors that could hinder understanding.
- **Completeness**: Determine if the predicted answer fully covers the scope of the ground truth answer. Does it leave out critical information or does it include all necessary details?
[…]
Explanation: <brief judgement of prediction>
Score: <a integer score of quality from 1-5>
@@
Your task is to act as an impartial and objective assessor of answers generated by a Large Multimodal Model (LMM) for video-based questions. Utilizing video frames, a posed question, and the models provided answer, your evaluation should focus on the following aspects:
[…]
- **Accuracy**: Compare the predicted answer to the ground truth answer. Does the prediction accurately reflect the information given in the ground truth answer without introducing factual inaccuracies?
- **Clarity**: Assess the clarity of the predicted answer. Look for issues such as repetition, unclear descriptions, or any grammatical errors that could hinder understanding.
- **Completeness**: Determine if the predicted answer fully covers the scope of the ground truth answer. Does it leave out critical information or does it include all necessary details?
[…]
Question: {question}
Model Predicted Answer: {prediction}
[…]
Explanation: <brief judgement of prediction>
Score: <an integer score of quality from 1-5>

### v2

@@ 4 Assessment of Evaluator with GPT-4V Caption as Evidence
FrozenBiLM (Yang et al., 2022)* 1B 32.2 - 16.8 - 41.0 -
VideoLLaMA (Zhang et al., 2023a)* 7B 51.6 2.5 29.6 1.8 - -
LLaMA-Adapter (Zhang et al., 2023b)* 7B 54.9 3.1 43.8 2.7 - -
VideoChat (Li et al., 2023b)* 7B 56.3 2.8 45.0 2.5 34.4 2.3
BT-Adapter Liu et al. (2023c)* 7B 67.5 3.7 57.0 3.2 - -
@@
Table 1: Evaluation of Model Performance on Zero-Shot Video Question Answering Benchmarks Using gpt-3.5-turbo-0613. Models denoted with * have their results directly sourced from their original publications. Caution is advised when interpreting these results; see Appendix A for an in-depth analysis of evaluation challenges. All other baseline models were reproduced by our team.
@@ Appendix F Prompts for GPT-4V and ChatGPT Queries
{minted}
[fontsize=, frame=single,linenos=false,breaklines,breaksymbol=,escapeinside=——,bgcolor=Box1Color]text Task Instructions:
Given a caption that summarizes the content of a video, generate three question-answer pairs that relate directly to the information and context provided in the caption. The questions should be grounded to the understanding of the video content.
Guidelines for QA Generation:
1. Helpfulness: Answers should provide sufficient detail and depth to fully address the question. They should include relevant explanations, or context where appropriate, to enhance understanding.
2. Faithfulness: The answers must accurately reflect the information presented in the video caption. Avoid speculation or the inclusion of information not contained or implied by the caption to maintain the integrity of the content.
3. Diversity: Craft questions that cover different aspects of the video caption to provide a comprehensive understanding of the content. This includes factual inquiries, inferential questions, and those that may elicit explanatory responses.
Input Video Caption: caption
Output format: Q1: ¡question1¿ A1: ¡answer1¿ Q2: ¡question2¿ A2: ¡answer2¿ Q3: ¡question3¿ A3: ¡answer3¿
@@
{minted}
[fontsize=, frame=single,linenos=false,breaklines,breaksymbol=,escapeinside=——,bgcolor=Box2Color]text Your role is to serve as an impartial and objective evaluator of a video caption provided by a Large Multimodal Model (LMM). Based on the input frames of a video, assess primarily on two criteria: the coverage of video elements in the caption and the absence of hallucinations in the response. In this context, ’hallucination’ refers to the model generating content not present or implied in the video, such as incorrect details about objects, actions, counts, or other aspects not evidenced in the video frames.
To evaluate the LMM’s response:
Start with a brief explanation of your evaluation process. Then, assign a rating from the following scale:
Rating 6: Very informative with good coverage, no hallucination Rating 5: Very informative, no hallucination Rating 4: Somewhat informative with some missing details, no hallucination Rating 3: Not informative, no hallucination Rating 2: Very informative, with hallucination Rating 1: Somewhat informative, with hallucination Rating 0: Not informative, with hallucination
LMM Response to Evaluate LLM_response
Output format: Judgment: ¡your judgment¿ Score: ¡integer value rating¿
@@
{minted}
[fontsize=, frame=single,linenos=false,breaklines,breaksymbol=,escapeinside=——,bgcolor=Box2Color]text Given the following inputs:
1. **Ground Truth Video Caption**: caption 2. **Question Related to the Caption**: question 3. **Ground Truth Answer**: answer 4. **Model Predicted Answer**: prediction
Your task is to evaluate the model’s predicted answer against the ground truth answer, based on the context provided by the video caption and the question. Consider the following criteria for evaluation:
- **Relevance**: Does the predicted answer directly address the question posed, considering the information provided in the video caption? - **Accuracy**: Compare the predicted answer to the ground truth answer. Does the prediction accurately reflect the information given in the ground truth answer without introducing factual inaccuracies? - **Clarity**: Assess the clarity of the predicted answer. Look for issues such as repetition, unclear descriptions, or any grammatical errors that could hinder understanding. - **Completeness**: Determine if the predicted answer fully covers the scope of the ground truth answer. Does it leave out critical information or does it include all necessary details?
**Output Format**: Explanation: ¡brief judgement of prediction¿ Score: ¡a integer score of quality from 1-5¿
@@
{minted}
[fontsize=, frame=single,linenos=false,breaklines,breaksymbol=,escapeinside=——,bgcolor=Box2Color]text Your task is to act as an impartial and objective assessor of answers generated by a Large Multimodal Model (LMM) for video-based questions. Utilizing video frames, a posed question, and the model’s provided answer, your evaluation should focus on the following aspects:
- **Relevance**: Does the predicted answer directly address the question posed, considering the information provided in the video caption? - **Accuracy**: Compare the predicted answer to the ground truth answer. Does the prediction accurately reflect the information given in the ground truth answer without introducing factual inaccuracies? - **Clarity**: Assess the clarity of the predicted answer. Look for issues such as repetition, unclear descriptions, or any grammatical errors that could hinder understanding. - **Completeness**: Determine if the predicted answer fully covers the scope of the ground truth answer. Does it leave out critical information or does it include all necessary details?
**Input**: Question: question Model Predicted Answer: prediction
**Output Format**: Explanation: ¡brief judgement of prediction¿ Score: ¡an integer score of quality from 1-5¿